# Final-Model-Only Data Attribution with a Unifying View of Gradient-Based Methods

### v1

@@ 4 Gradient-Based Methods as Approximate Further Training
The further training gold standard described in the previous section is computationally expensive. In the case of LOO further training, the number of further trainings scales with the training set size $n$, or at least the number of training instances for which we wish to estimate attribution scores. If we wish to approximate the expectation in (3) by averaging over multiple realizations of $\xi$, that further scales the cost. It is natural therefore to consider approximate further training. In this section, we assume that the amount of further training is limited, so that the resulting changes in model parameters, $\Delta{\bm{\theta}}$, are small. We show that several gradient-based methods for TDA can be re-derived from this perspective. This suggests that they are better viewed as approximations to further training and better suited to sensitivity analysis in the FiMO setting than for other purposes. It is important to note that we do not assume the given parameters ${\bm{\theta}}^{f}$ are a stationary point for training set $\mathcal{D}$.
@@ 4.1 Approximate further training
The assumption that $\Delta{\bm{\theta}}$ is small leads us to consider first- and second-order Taylor expansions of (1):
@@ 4.2 First-order methods
which is proportional to the inner product between training loss and test evaluation gradients. This corresponds to the “Grad-Dot” method (Charpiat et al., 2019), and also to a special case of TracIn (Pruthi et al., 2020) that uses only the final checkpoint. Here the damping value $\lambda$ turns out not to matter because it is an overall constant of proportionality. A variation is to use the cosine similarity between gradients (“Grad-Cos”) instead of the unnormalized inner product.
@@ 4.3.1 Influence functions for (4)
If ${\bm{\theta}}^{f}$ is not stationary however, then $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})\neq{\bm{0}}$ as discussed in Section 3. Previous works like Koh and Liang (2017) still neglect the last $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$ term in (7), arguing that if ${\bm{\theta}}^{f}$ is near-stationary, then $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$ is small and the product $\epsilon\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$ is second-order. It is also convenient to neglect this $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$ term because one would otherwise have to compute $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$. We will neglect it as well in the present work to be consistent with previous TDA methods. In Appendices B.2 and B.3 however, we provide an expression for $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$ and also show how to account for the $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$ term in (7) without requiring the Hessian $\nabla^{2}_{\bm{\theta}}L({\bm{z}}_{i};{\bm{\theta}}^{f})$.
@@ 4.3.2 Gauss-Newton approximate Hessian
It is often the case that the loss function $L({\bm{z}}_{i};{\bm{\theta}})$ is a composition ¯$\bar{\ell}(\bar{f}({\bm{z}}_{i};{\bm{\theta}}))$ of a scalar-valued model output function $\bar{f}({\bm{z}}_{i};{\bm{\theta}})$ with a convex univariate loss function ¯$\bar{\ell}(\bar{f})$. Here we allow $\bar{f}({\bm{z}}_{i};{\bm{\theta}})$ to also depend on the target $y_{i}$. For example, if $f({\bm{x}}_{i};{\bm{\theta}})$ is a regression model, $\bar{f}({\bm{z}}_{i};{\bm{\theta}})=f({\bm{x}}_{i};{\bm{\theta}})-y_{i}$ can be the prediction error. For multi-class classification, Park et al. (2023) define $\bar{f}({\bm{z}}_{i};{\bm{\theta}})$ to be the logit of the predicted probability of the target class $y_{i}$, which allows multi-class classification to be handled in the same way (see Section 3.3 of Park et al. (2023) for details).
Given the compositional form $L({\bm{z}}_{i};{\bm{\theta}})=\bar{\ell}(\bar{f}({\bm{z}}_{i};{\bm{\theta}}))$, the gradient and Hessian of $L({\bm{z}}_{i};{\bm{\theta}})$ are given by
$\nabla_{\bm{\theta}}L({\bm{z}}_{i};{\bm{\theta}})$ =¯$=\bar{\ell}^{\prime}(\bar{f}({\bm{z}}_{i};{\bm{\theta}}))\nabla_{\bm{\theta}}\bar{f}({\bm{z}}_{i};{\bm{\theta}}),$ (9)
$\nabla^{2}_{\bm{\theta}}L({\bm{z}}_{i};{\bm{\theta}})$ =¯′′$=\bar{\ell}^{\prime\prime}(\bar{f}({\bm{z}}_{i};{\bm{\theta}}))\nabla_{\bm{\theta}}\bar{f}({\bm{z}}_{i};{\bm{\theta}})\nabla_{\bm{\theta}}\bar{f}({\bm{z}}_{i};{\bm{\theta}})^{T}+\bar{\ell}^{\prime}(\bar{f}({\bm{z}}_{i};{\bm{\theta}}))\nabla^{2}_{\bm{\theta}}\bar{f}({\bm{z}}_{i};{\bm{\theta}}).$ (10)
@@
To obtain more compact expressions, let us define the gradient vectors ${\bm{g}}_{i}=\nabla_{\bm{\theta}}\bar{f}({\bm{z}}_{i};{\bm{\theta}}^{f})$, $i=1,\dots,n$ and matrix ${\bm{G}}=\begin{bmatrix}{\bm{g}}_{1}&\dots&{\bm{g}}_{n}\end{bmatrix}^{T}$, vector ${\bm{r}}=-\begin{bmatrix}\bar{\ell}^{\prime}(\bar{f}({\bm{z}}_{1};{\bm{\theta}}^{f}))&\dots&\bar{\ell}^{\prime}(\bar{f}({\bm{z}}_{n};{\bm{\theta}}^{f}))\end{bmatrix}^{T}$, and $n\times n$ diagonal matrix ${\bm{V}}$ with ¯′′$\bar{\ell}^{\prime\prime}(\bar{f}({\bm{z}}_{1};{\bm{\theta}}^{f})),\dots,\bar{\ell}^{\prime\prime}(\bar{f}({\bm{z}}_{n};{\bm{\theta}}^{f}))$ as its diagonal entries. Using these definitions, (9), and the Gauss-Newton simplification of (10), the quadratic optimization in (4) can be rewritten for $\mathcal{D}^{\prime}=\mathcal{D}$ as
@@
where we have also disregarded the constant term.
@@ 4.3.3 Relationships with existing influence function methods
Both (8) and (12) require computing an inverse Hessian-gradient product, with either the damped true Hessian ${\bm{H}}({\bm{\theta}}^{f})+\lambda{\bm{I}}$ or the Gauss-Newton Hessian ${\bm{G}}^{T}{\bm{V}}{\bm{G}}+\lambda{\bm{I}}$. For models with a large number of parameters $p$, the $O(p^{3})$ computational cost of directly solving the system of equations is prohibitive. Koh and Liang (2017) proposed to apply CG (Martens, 2010) and LiSSA (Agarwal et al., 2017), two methods that approximate the solution iteratively. The former does so by applying the CG algorithm to an equivalent quadratic minimization problem, while the latter uses a truncated Neumann series. Both use repeated Hessian-vector products (HVPs) and can be applied to either (8) or (12).
Arnoldi influence functions
Schioppa et al. (2022) propose to use Arnoldi iteration (a.k.a. Lanczos iteration for symmetric matrices) to approximately compute the largest-magnitude eigenvalues and corresponding eigenvectors of ${\bm{H}}({\bm{\theta}}^{f})$. The method then projects gradient vectors into the subspace of top eigenvectors of ${\bm{H}}({\bm{\theta}}^{f})$. By virtue of the orthonormal basis of eigenvectors for this subspace, the matrix inversion in (8) reduces to scalar division by eigenvalues.
@@
TRAK (Park et al., 2023) can be seen as a hybrid TDA method that combines a gradient-based approximation with re-training (training an ensemble of models). Under the FiMO setting where re-training is not possible, we focus on the gradient-based part of TRAK by setting its parameter $M$ (number of trained models) to $1$. This variant, which we refer to as TRAKM=1, falls under the Gauss-Newton framework of Section 4.3.2. First, TRAK applies a random projection matrix ${\bm{P}}\sim{\mathcal{N}}(0,1)^{p\times k}$ to reduce the dimensionality of the gradients: ${\bm{\phi}}_{i}={\bm{P}}^{T}{\bm{g}}_{i}$ and ${\bm{\Phi}}={\bm{G}}{\bm{P}}$. With ${\bm{\phi}}_{i}$ and ${\bm{\Phi}}$ in place of ${\bm{g}}_{i}$ and ${\bm{G}}$, TRAKM=1 corresponds to a special case of (12) with $\lambda=0$ (no regularization) and ${\bm{V}}={\bm{I}}$ (non-identity ${\bm{V}}$ found empirically to have little effect): [^3] Park et al. (2023) actually begin with an exact leave-one-out formula but find that the denominator also has little effect.
$\widehat{\Delta{\bm{\theta}}}_{\mathrm{TRAK}}(\mathcal{D}_{-i})=-r_{i}({\bm{\Phi}}^{T}{\bm{\Phi}})^{-1}{\bm{\phi}}_{i}.$
@@
Grosse et al. (2023) start with the Gauss-Newton approximation in their work, corresponding to (12). Next, they adopt the K-FAC and EK-FAC approximations from Martens and Grosse (2015); George et al. (2018), which first approximate the Gauss-Newton Hessian as block-diagonal, where each block corresponds to a layer in a NN. This allows influence functions to be computed separately for each layer. Then for certain layers, an uncorrelatedness assumption is made that permits a Kronecker factorization of the corresponding block of the Hessian, making inversion much more efficient.
@@
Kwon et al. (2024) start with a variant of the Gauss-Newton approximation in which ¯$\bar{\ell}(\bar{f})=\bar{f}$ is the identity function and ${\bm{g}}_{i}=\nabla_{\bm{\theta}}L({\bm{z}}_{i};{\bm{\theta}}^{f})$ is a gradient of the loss. They then make the same block-diagonal/layer-wise approximation as Grosse et al. (2023). Setting this simplification aside, DataInf corresponds to (12) with $r_{i}=-1$ (identity $\bar{\ell}$) and ${\bm{V}}=(1/n){\bm{I}}$ (identity $\bar{\ell}$, average over $\mathcal{D}$ instead of sum). The result is $\widehat{\Delta{\bm{\theta}}(\mathcal{D}_{-i})}=((1/n){\bm{G}}^{T}{\bm{G}}+\lambda{\bm{I}})^{-1}{\bm{g}}_{i}$. The key idea of DataInf is to then interchange the order of averaging and matrix inversion, i.e.,
$\left(\frac{1}{n}{\bm{G}}^{T}{\bm{G}}+\lambda{\bm{I}}\right)^{-1}=\left(\frac{1}{n}\sum_{i=1}^{n}{\bm{g}}_{i}{\bm{g}}_{i}^{T}+\lambda{\bm{I}}\right)^{-1}\approx\frac{1}{n}\sum_{i=1}^{n}\left({\bm{g}}_{i}{\bm{g}}_{i}^{T}+\lambda{\bm{I}}\right)^{-1}.$
They then use the Sherman-Morrison formula to express $({\bm{g}}_{i}{\bm{g}}_{i}^{T}+\lambda{\bm{I}})^{-1}$ in closed form, resulting in an approximation to (5) that involves only gradient inner products.
@@ 5 Related Work
This work builds upon existing works that also take a retrospective look at TDA methods, and influence functions in particular (Basu et al., 2021; Zhang and Zhang, 2022; Bae et al., 2022; Schioppa et al., 2023; Nguyen et al., 2023). We discuss individual papers more deeply in Appendix C. Here, we note the key differences between our work and these prior works:
@@
    More gradient-based methods: Our unified view (Section 4) and numerical comparisons (Section 6) encompass gradient-based methods beyond influence functions, including more recent methods such as TRAK and DataInf as well as first-order methods. We thereby obtain insights into similarities and differences among gradient-based methods, in addition to their performance on a task (approximation of further training in our case). In contrast, Basu et al. (2021); Zhang and Zhang (2022); Bae et al. (2022); Schioppa et al. (2023) mainly focus on influence functions, while Nguyen et al. (2023) evaluate fewer methods than we do.
@@
    Further training: The further training gold standard in Section 3 formalizes similar procedures used in previous works (in Appendix D.2 in Bae et al. (2022), in the experiment in Section 5.2 of Schioppa et al. (2023)). The proposal of averaging over realizations of stochastic further training appears to be new, and our results in Figure 3 show that it brings further training closer to what gradient-based methods estimate. Bae et al. (2022) propose an alternative gold standard called the proximal Bregman response function (PBRF). The PBRF however involves a non-standard Bregman distance objective, tailored specifically to be closer to influence functions, whereas our further training is more generic.
@@ 6 Numerical Comparison of Gradient-Based Methods to Further Training
In Section 4, we showed that many gradient-based TDA methods are approximations to further training in the FiMO setting. We now report on experiments that assess the quality of these approximations.
@@ 6.1 Experimental setup
Our experiments span the modalities of tabular, image, and text data. However, since we aimed to implement the further training gold standard described in Section 3, which is computationally expensive, our experiments are most comprehensive in the tabular case. We used four tabular datasets: two for regression, Concrete Strength and Energy Efficiency from the UCI repository (Kelly et al., ) following Bae et al. (2022), and two larger ones for classification, FICO Challenge (FICO, 2018) and Folktables (Ding et al., 2021). For image data, we chose the CIFAR-10 image classification dataset (Krizhevsky, 2009), while for text, we used the SST-2 sentiment classification dataset (Socher et al., 2013), which is part of the GLUE benchmark (Wang et al., 2019).
@@
For all tabular datasets, we used a 2-hidden-layer multi-layer perceptron (MLP) with 128 units in each hidden layer. The MLPs were trained using stochastic gradient descent (SGD) to yield the final model $f({\bm{x}},{\bm{\theta}}^{f})$. For CIFAR-10, we used a ResNet-9 architecture (He et al., 2015) and trained it using SGD. For SST-2, we fine-tuned a pre-trained BERT model (Devlin et al., 2019) using AdamW (Loshchilov and Hutter, 2019).
@@
Our aim was to estimate influence scores of a subset ${\mathcal{L}}\subset\mathcal{D}$ of training instances on the losses of a subset of $m$ test instances. Thus, the evaluation function was the loss, $g({\bm{z}},{\bm{\theta}})=L({\bm{z}};{\bm{\theta}})$. For the Concrete and Energy datasets, we considered all test instances, while for the other datasets, we selected $m=100$ test instances at random. The training subset ${\mathcal{L}}$ was also selected randomly with $l=\lvert{\mathcal{L}}\rvert=100$ for the tabular datasets and $l=50$ for CIFAR-10 and SST-2.
@@
For the results in this section, the further training algorithm $A^{\prime}$ is the same as $A$, i.e., AdamW for BERT and SGD otherwise. Appendix E has results on the tabular datasets using Adam as $A^{\prime}$ instead. We evaluated the approximation quality of gradient-based methods as a function of the amount of further training (epochs or steps). To approximate the expectation over $\xi$ in (3), we took averages over $r=100$ random seeds (denoted as realizations $\xi^{(s)}$ in (13) below), which control the order of instances in each training epoch.
We considered two methods for adjusting for the effect of further training alone. The first is as specified in (2), (3), i.e., subtracting the output due to further training on the full dataset $\mathcal{D}$. We found however that for some datasets, this adjustment left a non-negligible bias that varies from epoch to epoch, resulting in noisy cosine similarities. Please see Appendices D.3 and E for further discussion of and results from this first method. The second method (used in this section) subtracts the mean of the effects due to each LOO dataset $\mathcal{D}_{-i}$. Incorporating also the averaging over random seeds, the “gold” attribution score from further training is thus
@@
Approximate TDA methods
We experimented with the first-order methods Grad-Dot (Charpiat et al., 2019) (final-checkpoint-only TracIn (Pruthi et al., 2020)) and Grad-Cos, and influence function methods CG and LiSSA (Agarwal et al., 2017; Koh and Liang, 2017), TRAKM=1 (Park et al., 2023), EK-FAC (Grosse et al., 2023), and DataInf (Kwon et al., 2024). For LiSSA, we considered both the Gauss-Newton version ((12), referred to simply as “LiSSA”) as well as the true Hessian version ((8), LiSSA-H). For CG, only the Gauss-Newton version was used, while Gauss-Newton is the only choice for TRAKM=1, EK-FAC, and DataInf. CG could not be run on CIFAR-10 and SST-2 because it took too long, whereas LiSSA-H ran out of GPU memory on SST-2. We adapted TRAK for regression (see Appendix D.4).
@@
For each test instance evaluated, we have a vector ${\bm{a}}\in{}^{l}$ of gold attribution scores from further training (13), and corresponding vectors $\hat{{\bm{a}}}$ from gradient-based methods. We use the cosine similarity between ${\bm{a}}$ and $\hat{{\bm{a}}}$ as the evaluation metric.

### v2

@@ 4 Gradient-Based Methods as Approximate Further Training
The further training gold standard described in the previous section is computationally expensive. In the case of LOO further training, the number of further trainings scales with the training set size $n$, or at least the number of training instances for which we wish to estimate attribution scores. If we wish to approximate the expectation in (3) by averaging over multiple realizations of $\xi$, that further scales the cost. It is natural therefore to consider approximate further training. In this section, we assume that the amount of further training is limited, so that the resulting changes in model parameters, $\Delta{\bm{\theta}}$, are small. We show that several gradient-based methods for TDA can be re-derived from this perspective. We argue therefore that they are better viewed as approximations to further training and better suited to FiMODA than to other TDA variants.
It is important to note that unlike the standard influence function literature, we do not assume that the empirical risk $R(\mathcal{D}^{\prime},{\bm{\theta}})$ is convex, nor do we assume that the given parameters ${\bm{\theta}}^{f}$ are a stationary point of $R(\mathcal{D},{\bm{\theta}})$ over the full training set $\mathcal{D}$. As a result of the latter, we obtain generalized influence function expressions (Proposition 1 and Corollary 2). While we then specialize to the standard expressions used in existing works, the generalized expressions could be useful in future work.
@@ 4.1 Approximate further training
The assumption of small $\Delta{\bm{\theta}}$ leads us to consider first- and second-order Taylor expansions of (1):
@@ 4.2 First-order methods
proportional to the inner product between training loss and test evaluation gradients. This corresponds to the “Grad-Dot” method (charpiat2019similarity), and to a special case of TracIn (pruthi2020tracin) that uses only the final checkpoint. A variation is to use cosine similarity (“Grad-Cos”) instead of the unnormalized inner product.
@@ 4.3.1 Influence functions for (4)
If ${\bm{\theta}}^{f}$ is not stationary however, then $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})\neq{\bm{0}}$ as discussed in Section 3. Previous works koh2017inffunc neglect the last $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$ term in (7), arguing that if ${\bm{\theta}}^{f}$ is near-stationary, then $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$ is small and the product $\epsilon\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$ is second-order. It is also convenient to neglect this term to avoid computing $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$. We will neglect it as well in the main paper to be consistent with previous TDA methods. In Appendices B.2 and B.3 however, we provide an expression for $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$ and also show how to account for the $\widehat{\Delta{\bm{\theta}}}(\mathcal{D})$ term in (7) without requiring the Hessian $\nabla^{2}_{\bm{\theta}}L({\bm{z}}_{i};{\bm{\theta}}^{f})$.
@@ 4.3.2 Gauss-Newton approximate Hessian
It is often the case that the loss function $L({\bm{z}}_{i};{\bm{\theta}})$ is a composition $\bar{\ell}(\bar{f}({\bm{z}}_{i};{\bm{\theta}}))$ of a scalar-valued model output function $\bar{f}({\bm{z}}_{i};{\bm{\theta}})$ with a convex univariate loss function $\bar{\ell}(\bar{f})$. Here we allow $\bar{f}({\bm{z}}_{i};{\bm{\theta}})$ to also depend on the target $y_{i}$. For example, if $f({\bm{x}}_{i};{\bm{\theta}})$ is a regression model, $\bar{f}({\bm{z}}_{i};{\bm{\theta}})=f({\bm{x}}_{i};{\bm{\theta}})-y_{i}$ can be the prediction error. For multi-class classification, park2023trak define $\bar{f}({\bm{z}}_{i};{\bm{\theta}})$ to be the logit of the predicted probability of the target class $y_{i}$, which allows multi-class classification to be handled in the same way (see (park2023trak, Sec. 3.3) for details).
Given the composition $L({\bm{z}}_{i};{\bm{\theta}})=\bar{\ell}(\bar{f}({\bm{z}}_{i};{\bm{\theta}}))$, the gradient and Hessian of $L({\bm{z}}_{i};{\bm{\theta}})$ are given by
$\nabla_{\bm{\theta}}L({\bm{z}}_{i};{\bm{\theta}})$ $=\bar{\ell}^{\prime}(\bar{f}({\bm{z}}_{i};{\bm{\theta}}))\nabla_{\bm{\theta}}\bar{f}({\bm{z}}_{i};{\bm{\theta}}),$ (9)
$\nabla^{2}_{\bm{\theta}}L({\bm{z}}_{i};{\bm{\theta}})$ $=\bar{\ell}^{\prime\prime}(\bar{f}({\bm{z}}_{i};{\bm{\theta}}))\nabla_{\bm{\theta}}\bar{f}({\bm{z}}_{i};{\bm{\theta}})\nabla_{\bm{\theta}}\bar{f}({\bm{z}}_{i};{\bm{\theta}})^{T}+\bar{\ell}^{\prime}(\bar{f}({\bm{z}}_{i};{\bm{\theta}}))\nabla^{2}_{\bm{\theta}}\bar{f}({\bm{z}}_{i};{\bm{\theta}}).$ (10)
@@
To obtain more compact expressions, let us define the gradient vectors ${\bm{g}}_{i}=\nabla_{\bm{\theta}}\bar{f}({\bm{z}}_{i};{\bm{\theta}}^{f})$, $i=1,\dots,n$ and matrix ${\bm{G}}=\begin{bmatrix}{\bm{g}}_{1}&\dots&{\bm{g}}_{n}\end{bmatrix}^{T}$, vector ${\bm{r}}=-\begin{bmatrix}\bar{\ell}^{\prime}(\bar{f}({\bm{z}}_{1};{\bm{\theta}}^{f}))&\dots&\bar{\ell}^{\prime}(\bar{f}({\bm{z}}_{n};{\bm{\theta}}^{f}))\end{bmatrix}^{T}$, and $n\times n$ diagonal matrix $\bm{V}$} with $\bar{\ell}^{\prime\prime}(\bar{f}({\bm{z}}_{1};{\bm{\theta}}^{f})),\dots,\bar{\ell}^{\prime\prime}(\bar{f}({\bm{z}}_{n};{\bm{\theta}}^{f}))$ as its diagonal entries. Using these definitions, (9), and the Gauss-Newton simplification of (10), and disregarding the constant term, the quadratic optimization in (4) can be rewritten for $\mathcal{D}^{\prime}=\mathcal{D}$ as
@@ 4.3.3 Relationships with existing influence function methods
Both (8) and (12) require computing an inverse Hessian-gradient product, with either the damped true Hessian ${\bm{H}}({\bm{\theta}}^{f})+\lambda{\bm{I}}$ or the Gauss-Newton Hessian ${\bm{G}}^{T}{\bm{V}}{\bm{G}}+\lambda{\bm{I}}$. For models with a large number of parameters $p$, the $O(p^{3})$ computational cost of directly solving the system of equations is prohibitive. koh2017inffunc proposed to apply CG (martens2010deep) and LiSSA (agarwal2017second), two methods that approximate the solution iteratively. The former does so by applying the CG algorithm to an equivalent quadratic minimization problem, while the latter uses a truncated Neumann series. Both can be applied to either (8) or (12).
@@
TRAK (park2023trak) can be seen as a hybrid TDA method that combines a gradient-based approximation with re-training (training an ensemble of models). For FiMODA where re-training is not possible, we focus on the gradient-based part of TRAK by setting its parameter $M$ (number of trained models) to $1$. This variant, which we refer to as TRAKM=1, falls under the Gauss-Newton framework of Section 4.3.2. First, TRAK applies a random projection matrix ${\bm{P}}\sim{\mathcal{N}}(0,1)^{p\times k}$ to reduce the dimensionality of the gradients: ${\bm{\phi}}_{i}={\bm{P}}^{T}{\bm{g}}_{i}$ and ${\bm{\Phi}}={\bm{G}}{\bm{P}}$. With ${\bm{\phi}}_{i}$ and $\bm{\Phi}$} in place of ${\bm{g}}_{i}$ and $\bm{G}$}, TRAKM=1 corresponds to a special case of (12) with $\lambda=0$ (no regularization) and ${\bm{V}}={\bm{I}}$ (non-identity $\bm{V}$} found empirically to have little effect), i.e., $\widehat{\Delta{\bm{\theta}}}_{\mathrm{TRAK}}(\mathcal{D}_{-i})=-r_{i}({\bm{\Phi}}^{T}{\bm{\Phi}})^{-1}{\bm{\phi}}_{i}.$
@@
grosse2023studyingllmgen start with the Gauss-Newton approximation in their work, corresponding to (12). Next, they adopt the K-FAC and EK-FAC approximations from martens2015kfac; george2018ekfac, which first approximate the Gauss-Newton Hessian as block-diagonal, where each block corresponds to a layer in a NN. This allows influence functions to be computed separately for each layer. Then for certain layers, an uncorrelatedness assumption is made that permits a Kronecker factorization of the corresponding block of the Hessian, making inversion much more efficient.
@@
kwon2024datainf start with a variant of the Gauss-Newton approximation in which $\bar{\ell}(\bar{f})=\bar{f}$ is the identity function and ${\bm{g}}_{i}=\nabla_{\bm{\theta}}L({\bm{z}}_{i};{\bm{\theta}}^{f})$ is a gradient of the loss. They then make the same block-diagonal/layer-wise approximation as grosse2023studyingllmgen. Setting this simplification aside, DataInf corresponds to (12) with $r_{i}=-1$ (identity $\bar{\ell}$) and ${\bm{V}}=(1/n){\bm{I}}$ (identity $\bar{\ell}$, average over $\mathcal{D}$ instead of sum). The result is $\widehat{\Delta{\bm{\theta}}(\mathcal{D}_{-i})}=((1/n){\bm{G}}^{T}{\bm{G}}+\lambda{\bm{I}})^{-1}{\bm{g}}_{i}$. The key idea of DataInf is to then interchange the order of averaging and matrix inversion (please see kwon2024datainf).
@@ 5 Related Work
This work builds upon existing works that also take a retrospective look at TDA methods, and influence functions especially (basu2021inffragile; zhang2022rethinking; bae2022infanswer; schioppa2023theoretical; nguyen2023bayesian). We discuss individual papers more in Appendix C. Here, we note key differences between our work and these prior works (summarized in Table 1):
@@
    More gradient-based methods: Our unified view (Section 4) and numerical comparisons (Section 6) encompass gradient-based methods beyond influence functions, including more recent ones such as TRAK and DataInf as well as first-order methods. We thereby obtain insights into similarities and differences among gradient-based methods, in addition to how well they approximate further training. In contrast, basu2021inffragile; zhang2022rethinking; bae2022infanswer; schioppa2023theoretical mainly focus on influence functions, while nguyen2023bayesian evaluate fewer methods than we do.
@@
    Further training: The further training gold standard in Section 3 formalizes and refines similar procedures used in previous works (in Appendix D.2 in bae2022infanswer, in Section 5.2 of schioppa2023theoretical). The proposal of averaging over realizations of stochastic further training appears to be new, and our results in Figures 3 and 7 show that it brings further training closer to what gradient-based methods estimate. bae2022infanswer propose an alternative gold standard called the proximal Bregman response function (PBRF). The PBRF however involves a non-standard Bregman distance objective, tailored specifically to be closer to influence functions, whereas our further training is more generic.
@@ 6 Numerical Comparison of Gradient-Based Methods to Further Training
In Section 4, we showed that many gradient-based TDA methods are approximations to further training for FiMODA. We now report on experiments that assess the quality of these approximations. Code to help reproduce these experiments is provided at https://github.com/IBM/fimoda.
@@ 6.1 Experimental setup
Our experiments span the modalities of tabular, image, and text data. However, since we aimed to implement the further training gold standard described in Section 3, which is computationally expensive, our experiments are most comprehensive in the tabular case. We used four tabular datasets: two for regression, Concrete Strength and Energy Efficiency from the UCI repository (uci) following bae2022infanswer, and two larger ones for classification, FICO Challenge (fico2018explainable) and Folktables (ding2021retiring). For image data, we chose the CIFAR-10 image classification dataset (cifar), while for text, we used the SST-2 sentiment classification dataset (sst2), which is part of the GLUE benchmark (glue).
@@
For all tabular datasets, we used a 2-hidden-layer multi-layer perceptron (MLP) with 128 units in each hidden layer. The MLPs were trained using stochastic gradient descent (SGD) to yield the final model $f({\bm{x}},{\bm{\theta}}^{f})$. For CIFAR-10, we used a ResNet-9 architecture (resnet) and trained it using SGD. For SST-2, we fine-tuned a pre-trained BERT model (bert) using AdamW (loshchilov2018decoupled).
@@
Our aim was to estimate influence scores of a subset ${\mathcal{L}}\subset\mathcal{D}$ of training instances on the losses of a subset of $m$ test instances. Thus, the evaluation function was the loss, $g({\bm{z}},{\bm{\theta}})=L({\bm{z}};{\bm{\theta}})$. For the Concrete and Energy datasets, we considered all test instances, while for the other datasets, we selected $m=100$ test instances at random. The training subset $\mathcal{L}$} was also selected randomly with $l=\lvert{\mathcal{L}}\rvert=100$ for the tabular datasets and $l=50$ for CIFAR-10 and SST-2. The training subset size $l$ was limited by our aim to conduct further training, since the number of further trainings is proportional to $l$. Nevertheless, $l=50$ or $100$ is greater than the number of training instances used in similar retrospective work on influence functions (bae2022infanswer used 20 instances according to their Appendix C.2, schioppa2023theoretical used 32 in their Section 5.2).
@@
For the results in this section, the further training algorithm $A^{\prime}$ is the same as $A$, i.e., AdamW for BERT and SGD otherwise. Appendix E.1 has results on the tabular datasets using Adam as $A^{\prime}$ instead. We evaluated the approximation quality of gradient-based methods as a function of the amount of further training (epochs or steps). To approximate the expectation over $\xi$ in (3), we took averages over $r=100$ random seeds (denoted as realizations $\xi^{(s)}$ in (13) below), which control the order of instances in each training epoch. To achieve the effect of leaving out a training instance ${\bm{z}}_{i}$, we followed schioppa2023theoretical in subtracting a second loss term $L({\bm{z}}_{i};{\bm{\theta}})/n$ from the usual loss for every batch (please see Appendix D.3 for explanation of the $1/n$ scaling). While schioppa2023theoretical did not provide a justification for this practice, we believe that it also mitigates the randomness of stochastic mini-batch training, since the mini-batch that ${\bm{z}}_{i}$ appears in/is left out of is random.
We considered two methods for adjusting for the effect of further training alone. The first is as specified in (2), (3), i.e., subtracting the output due to further training on the full dataset $\mathcal{D}$. We found however that for some datasets, this adjustment left a non-negligible bias that varies from epoch to epoch, resulting in noisy cosine similarities. Please see Appendices D.3 and E.1 for further discussion of and results from this first method. The second method (used in this section) subtracts the mean of the effects due to each LOO dataset $\mathcal{D}_{-i}$. Incorporating also the averaging over random seeds, the “gold” attribution score from further training is thus
@@
Approximate FiMODA methods
We experimented with the first-order methods Grad-Dot (charpiat2019similarity) (final-checkpoint-only TracIn (pruthi2020tracin)) and Grad-Cos, and influence function methods CG and LiSSA (agarwal2017second; koh2017inffunc), TRAKM=1 (park2023trak), EK-FAC (grosse2023studyingllmgen), and DataInf (kwon2024datainf). For LiSSA, we considered both the Gauss-Newton version ((12), referred to simply as “LiSSA”) as well as the true Hessian version ((8), LiSSA-H). For CG, only the Gauss-Newton version was used. CG could not be run on CIFAR-10 and SST-2 because it took too long, whereas LiSSA-H ran out of GPU memory on SST-2. For both CG and LiSSA, we used a damping value of $\lambda=0.01$, which we found to yield a better approximation to further training than the $\lambda=0.001$ value used by bae2022infanswer (see Appendix D.4 for other parameter settings). We adapted TRAK for regression (see Appendix D.4).
@@
For each test instance evaluated, we have a vector ${\bm{a}}\in\real^{l}$ of gold attribution scores from further training (13), and corresponding vectors $\hat{{\bm{a}}}$ from gradient-based methods. We use the cosine similarity between $\bm{a}$} and $\hat{{\bm{a}}}$ as the evaluation metric. The choice of cosine similarity is motivated by the following reasons elaborated on in Appendix D.5: 1) It measures similarity in signed magnitudes between $\bm{a}$} and $\hat{{\bm{a}}}$, which is in line with the derivations in Section 4; 2) it is not sensitive however to an overall scale difference between $\bm{a}$} and $\hat{{\bm{a}}}$; 3) it is a more demanding measure than Pearson correlation. Appendix E.1 presents qualitatively similar results using Spearman rank correlation, a common metric in the TDA literature.